OSMOS Hardware Control Failure

Monday, April 8, 2013 2:56 PM

Problem(s) Encountered:

Experienced problems with Filter1 communicating with the IP.

Disperser, Slit, and Filter2 wheels all worked fine.


The normal restart procedure for "stuck" mechanisms in OSMOS is

   0) Leave the IC and Head Electronics Box alone (they have play no role with the mechanisms)

   1) Quit out of the osmos IE agent program (mdmTools stop osmos)

   2) Power cycle the IE

   3) Restart the OSMOS IE agent program (mdmTools start osmos)

   4) Type "startup" in prospero

   5) Run ossetup in prospero

The last step does a full reset of all the mechanisms after the power cycling. (Rick Pogge)